# 3AM: 3egment Anything with Geometric Consistency in Videos

### v3

@@ Abstract
Video object segmentation methods like SAM2 achieve strong performance through memory-based architectures but struggle under large viewpoint changes due to reliance on appearance features. Traditional 3D instance segmentation methods address viewpoint consistency but require camera poses, depth maps, and expensive preprocessing. We introduce 3AM, a training-time enhancement that integrates 3D-aware features from MUSt3R into SAM2. Our lightweight Feature Merger fuses multi-level MUSt3R features that encode implicit geometric correspondence. Combined with SAM2’s appearance features, the model achieves geometry-consistent recognition grounded in both spatial position and visual similarity. We propose a field-of-view aware sampling strategy ensuring frames observe spatially consistent object regions for reliable 3D correspondence learning. Critically, our method requires only RGB input at inference, with no camera poses or preprocessing. On challenging datasets with wide-baseline motion (ScanNet++, Replica), 3AM substantially outperforms SAM2 and extensions, achieving 90.6% IoU and 71.7% Positive IoU on ScanNet++’s Selected Subset, improving over state-of-the-art VOS methods by +15.9 and +30.4 points. Project page: https://jayisaking.github.io/3AM-Page/
@@ 1 Introduction
Both paradigms have critical gaps (Fig. 2). 2D methods like SAM2 excel at efficiency but lack geometric awareness, failing under viewpoint variation (Fig. 2a). MOSEv2 [14] shows significant degradation under wide-baseline changes. Conversely, 3D methods achieve better consistency but require explicit 3D inputs (Fig. 2b) and struggle to generalize. 2D-to-3D lifting suffers from view-inconsistent predictions [23, 42, 53, 70, 84]. End-to-end approaches like PanSt3R [85] require offline processing. Can we achieve 3D-aware, viewpoint-robust segmentation while preserving promptability and efficiency, without explicit 3D supervision? (Fig. 2c)
We introduce 3AM (Fig. 1), a training-time enhancement to SAM2 that integrates 3D-aware features from reconstruction models. Our key insight is that MUSt3R [6] encodes implicit geometric correspondence through features learned from multi-view consistency. By fusing multi-level MUSt3R features with SAM2’s appearance features via a lightweight Feature Merger, we enable geometry-consistent recognition based on spatial position. This geometric awareness is learned from posed RGB video during training, but at inference requires only RGB input and prompts. We introduce a field-of-view aware sampling strategy ensuring frames observe spatially consistent object regions. On ScanNet++ [82] and Replica [50] with wide-baseline motion, 3AM achieves 90.6% IoU and 71.7% Positive IoU versus SAM2Long’s 74.7% and 41.3% (+15.9 and +30.4 points).
@@
  • Comprehensive experimental validation demonstrating substantial improvements over SAM2 on challenging datasets (ScanNet++, Replica), particularly on object reappearance and viewpoint changes.
@@ 2 Related Work
2D Video Object Segmentation.
Video object segmentation (VOS) segments and tracks target objects across video sequences with temporally coherent masks [43, 7, 57, 37, 64]. Early approaches relied on appearance propagation or optical flow but struggled with long-term consistency and occlusion [74, 40, 3]. Recent memory-based architectures leverage spatio-temporal attention and feature retrieval for improved object association [78, 10, 11, 35, 41, 69, 79, 77, 31, 12]. SAM2 [46], a promptable model with streaming memory supporting points, boxes, and masks, achieves strong results across benchmarks. Follow-up efforts extend SAM2 [72, 75, 81, 15, 56]: SAM2Long [15] introduces memory trees for long-range identity preservation, while DAM4SAM [56] incorporates distractor-aware updates for cluttered scenes. Despite these advances, large viewpoint changes, occlusion, and disappearance-reappearance events cause significant degradation, as shown by MOSEv2 [14, 68, 45]. To address these limitations, our method integrates 3D-aware representations into SAM2 for object-consistent segmentation under challenging viewpoint and appearance variations.
3D Instance Segmentation.
Mainstream 3D instance segmentation follows two tracks. Proposal-centric methods operate natively in 3D, predicting instance masks from point clouds with learned proposals, often requiring large-scale 3D supervision and suffering from domain shifts [48, 24, 18, 59, 38, 52, 28, 53, 44]. A complementary line lifts 2D results to 3D by projecting per-view masks and merging across views, but still depends on posed RGB-D and point clouds for mask consolidation [76, 70, 39, 4, 26, 66, 27, 58, 25, 2, 20, 55, 32]. Pure 2D promptable segmentation rarely suffices for object-aligned 3D instances because cross-view association and temporal consistency are not enforced by image-space inference alone. Consequently, multi-view fusion often exhibits view-inconsistent masks, error accumulation, and fragmented 3D instances [84, 70, 42, 53, 23]. In contrast, our approach imposes 3D-aware self-consistency directly on video, preserving identity and geometry without 3D annotations or point-cloud mask merging.
@@
End-to-End 3D-Aware Methods.
Recent progress in end-to-end 3D reconstruction shows modern models can directly infer geometric structure from 2D inputs, enhancing downstream 3D perception tasks [6, 63, 62, 61, 73, 9, 29, 65, 8, 80, 60, 33, 49, 16, 36, 51]. This has fueled end-to-end 3D-aware segmentation frameworks that unify geometric reasoning and object parsing [21, 22, 30]. PanSt3R [85] integrates 3D-aware features from MUSt3R [6] with a transformer decoder to jointly predict geometry and panoptic segmentation, achieving multi-view consistent segmentation through learnable quadratic fusion. However, it is incompatible with promptable segmentation backbones (e.g., SAM2), limiting interactive inputs and open-vocabulary segmentation. Moreover, it requires offline access to entire image sequences, preventing online or streaming use. Our method introduces 3D-aware features into the SAM2 framework, pairing geometric consistency with promptability and broad generalization to enable object-consistent segmentation and tracking without additional 3D supervision.
@@ 3.1 Problem Setting
Given an RGB image $\mathbf{I}$ and a user-provided prompt $p$ (which can be a mask, box, or point), our goal is to identify the same object throughout a video sequence while maintaining consistent segmentation, i.e., avoiding false correspondences or identity switches across frames. Previous approaches often fail to achieve this goal. 2D-based methods (e.g., SAM2[46]) rely solely on image-space appearance features, and thus are sensitive to camera viewpoint or distance changes. They also struggle with visually similar or confusing objects, as partially addressed by DAM4SAM[56], leading to frequent tracking failures as the frame sequence grows longer. 3D-based methods (e.g., Mask3D[48], OpenYOLO3D[4], EmbodiedSAM[71]) attempt to localize objects in space using 3D mask proposals or dense 2D masks combined with camera poses, but this requires running structure-from-motion (SfM) or similar pipelines. As the number of video frames increases, both runtime and computational complexity grow rapidly, making these approaches impractical for long sequences or large scenes. To address these limitations, we aim to exploit 3D-aware information while maintaining a 2D segmentation framework. Specifically, we encourage the model to perform feature matching that is geometry-consistent, enabling it to identify the same object by recognizing its correspondence in 3D space.
@@ 3.2 Architecture
The overall pipeline is shown in Fig. 3. For each video frame $i$, our model extracts two complementary feature streams that capture appearance cues and multi-view geometric consistency. First, the frame is processed by the SAM2 vision encoder, producing a 2D appearance feature map $F_{\mathrm{2D}}^{(i)}$. In parallel, the same frame is passed into MUSt3R, which attends to its own multi-view memory through MUSt3R’s internal cross-attention mechanism, yielding a geometry-aware feature map $F_{{3D}}^{i}$. Next, $F_{{2D}}^{(i)}$ and $F_{{2D}}^{i}$ are fused into a refined representation $F_{{merged}}^{i}$. The merged feature is processed by the memory attention module and the mask decoder, and is finally encoded by the memory encoder to produce memory features that will be referenced by future frames. We provide comparison of different memory selection methods in Sec. 4.
Feature Merging.
@@
As shown in Fig. 4, SAM2 becomes unreliable when the object reappears or when the viewpoint changes drastically. The heat map in the lower row demonstrates that SAM2 cannot reliably associate the query point with its target under large angular differences, revealing its limited capacity to maintain object identity when the visual appearance changes substantially. To address this issue, we enhance SAM2 with MUSt3R features. The upper row of Fig. 4 visualizes MUSt3R activations across depth. Early layers preserve a clearer correspondence pattern with the query point, while deeper layers produce more diffuse responses. This shift reflects MUSt3R’s progression toward point-cloud decoding: the representation becomes more geometry-oriented but less semantically aligned with the 2D query. Consequently, relying solely on either early or late layers is insufficient; early layers lack explicit geometric structure, whereas late layers lose semantic clarity. We therefore sample MUSt3R features from both early and late stages to capture complementary information. These multi-level MUSt3R features are processed by our Feature Merger, which jointly performs cross-attention fusion and convolutional refinement. The cross-attention blocks first integrate the sampled MUSt3R layers into a single geometry-aware feature, aligning information from different depths. Specifically, as shown in Fig. 4, we take the MUSt3R encoder feature at the shallowest sampled depth, i.e., the most semantically rich layer, as the initial input to the Feature Merger. It first passes through a self-attention block to establish an initial semantic representation. The remaining sampled MUSt3R features are then incorporated one by one through a sequence of cross-attention layers, where the current merged representation acts as the query and each additional MUSt3R feature provides the key–value pairs. In this way, the Feature Merger progressively integrates information from both shallow and deeper stages of MUSt3R. The output is then fed into a convolutional stage that merges it with SAM2’s 2D feature $F_{2D}$ for restoring fine spatial detail. We further compare different 3D foundation models in the ablation study and more architecture details and comparison are provided in the supplementary materials. The final merged feature, $F_{{merged}}$, retains SAM2’s high-resolution, segmentation-aware appearance cues while incorporating MUSt3R’s geometry-aware consistency, enabling robust object localization even under severe viewpoint changes.
Mask Decoding.
The output of the memory-attended features is fed into the SAM2 mask decoder to produce the segmentation mask for frame $i$. The resulting merged features are then encoded and appended to the memory, enabling temporally consistent object identification in subsequent frames.
@@ 3.3 Training Frame Sampling
Naïve Random Sampling.
A straightforward strategy is to randomly sample $N$ frames from the video as in Fig. 5(b). This exposes the model to a broad distribution of camera viewpoints and object appearances. Random sampling also naturally introduces distracting or visually similar objects in the scene , which encourages the model to rely on 3D-aware alignment rather than purely local 2D similarity.
Object Spanning Problem
@@
Field-of-View Sampling.
@@ 4 Experiments
To mitigate the imperfect 2D mask annotations common in 3D-based datasets, we train on a hybrid mixture of synthetic, real, and video segmentation datasets. Specifically, the Aria Synthetic Dataset(ASE) [1] provides clean geometric supervision, ScanNet++ [82] offers realistic indoor 3D consistency, and MOSE[13] maintains the original temporally coherent masks. This combination enables balanced spatial and temporal learning. Note that to remain consistent with our sampling strategy described in Sec. 3.3, we apply continuous sampling on MOSE, as it does not provide camera poses or depth maps. For ASE and ScanNet++, which include calibrated geometry, we adopt a mixed policy: continuous sampling with probability 0.2 and our FOV-aware sampling with probability 0.8. The FOV threshold $\tau$ is set to 0.25. To apply our sampling strategy, we first run MUSt3R inference on every scene in ScanNet++ and MOSE and store the resulting features. These precomputed features are then used during training to enable FOV-aware sampling. For continuous sampled batches, MUSt3R is executed on the fly.
@@
To address this limitation, we use 3D-based datasets, ScanNet++ [82] and Replica [50], which naturally provide extensive camera trajectory variation due to their 3D reconstruction requirements. Their wide-baseline viewpoints make them suitable for evaluating consistency under severe pose changes.
For each scene in both datasets, we use ground-truth camera poses to project 3D instance masks onto the 2D frames. We then select the frame with the largest visible mask area as the conditioning frame and perform forward and backward tracking from that reference.
Compared Methods.
For comparison, we use SAM2 as the state-of-the-art VOS baseline. We also include the following works: SAM2Long [15] and DAM4SAM [56], which enhance SAM2 by improving its memory selection for more robust tracking. For a fair comparison, we additionally finetune the original SAM2 under our dataset setting. Note that our method, 3AM, uses the original SAM2 memory-selection mechanism in all reported results unless specified otherwise.
Metrics.
We use three complementary metrics to evaluate tracking quality. IoU is computed over all frames, including those where the object is absent, reflecting overall stability. Positive IoU measures accuracy only on frames where a ground-truth mask is present. Successful IoU further restricts to visible frames where the prediction has non-zero overlap with ground truth, capturing accuracy conditioned on successful localization. These metrics assess robustness to disappearance and accuracy when tracking succeeds.
4.1.1 ScanNet++
Table 2: Video Object Segmentation results on the ScanNet++ [82] dataset. Positive IoU is computed over frames with ground-truth masks. Successful IoU is computed over frames where a ground-truth object exists and the IoU of prediction and GT $\neq 0$.
Method Whole Set Selected Subset
IoU $\uparrow$ Positive IoU $\uparrow$ Successful IoU $\uparrow$ IoU $\uparrow$ Positive IoU $\uparrow$ Successful IoU $\uparrow$
SAM2 [46] 0.4392 0.0235 0.0831 0.3397 0.0179 0.0395
SAM2 (Finetuned) 0.7815 0.0106 0.0657 0.6628 0.0038 0.0276
SAM2Long [15] 0.8233 0.4166 0.6855 0.7474 0.4133 0.6382
DAM4SAM [56] 0.8205 0.4193 0.6783 0.7648 0.4356 0.6650
3AM (Ours) 0.8898 0.5630 0.7155 0.9061 0.7168 0.7737
@@ 4.1.1 ScanNet++
The results are shown in Tab. 2, and we provide visualization at Fig. 8. On the ScanNet++ dataset, we further evaluate performance on a Selected Subset specifically constructed to emphasize object reappearance. For each object track, we count the number of non-continuous visible segments, where a segment is defined as any contiguous span of visible frames whose length $\ell$ exceeds a minimal threshold $\ell_{\min}$. Objects with frequent reappearance are typically associated with large viewpoint changes, since each disappearance–reappearance cycle often corresponds to observing the object from a substantially different angle or position. This subset therefore provides a focused evaluation of robustness under severe pose variation.
On the Whole Set, 3AM achieves the highest performance across all metrics, with an IoU of 0.8898, outperforming SAM2 (0.4392), SAM2Long (0.8233), and DAM4SAM (0.8205). 3AM also attains the best Positive IoU (0.5630) and Successful IoU (0.7155).
On the more challenging Selected Subset, 3AM maintains a clear advantage. It reaches an IoU of 0.9061 and a Positive IoU of 0.7168, substantially higher than SAM2Long (0.7474 / 0.4133) and DAM4SAM (0.7648 / 0.4356). 3AM achieves the best Successful IoU at 0.7737, highlighting its ability to maintain object identity even when the object reappears under large changes in viewpoint and spatial configuration.
One thing to note is that SAM2 performs worse after finetuning. We attribute this degradation to the absence of grounding information from MUSt3R, which causes the memory attention to be mistrained due to insufficient reference cues for distinguishing different objects.
Two-view Matching Comparison with SegMASt3R.
We further compare Our 3AM with SegMASt3R [22], a two-view matching approach built on MASt3R that associates segment masks between two images under large viewpoint changes. SegMASt3R is not directly prompable and does not natively support >2>2 multiview tracking. Nevertheless, we include it as a strong two-view baseline by adopting the following protocol: we use the ground-truth initial mask as the source-frame candidate, and for each subsequent frame we independently perform two-view mask matching against the source frame. To fairly compare the two approaches, when tracking with our 3AM, we don’t use any memory other than the source frame.
Under this protocol, 3AM achieves Reference ALL IoU of $0.8915$, POS IoU of $0.5115$, and SUC IoU of $0.6405$, whereas SegMASt3R obtains $0.6800$, $0.3628$, and $0.4053$, respectively. These results demonstrate 3AM not only provides stronger functionality, but also outperforms SegMASt3R even when restricted to two-view matching.
Table 3: Quantitative comparison between 3AM and SegMASt3R under the two-view matching evaluation protocol on ScanNet++ Whole Set.
Method IoU $\uparrow$ Positive IoU $\uparrow$ Successful IoU $\uparrow$
@@
3AM (Ours) 0.8915 0.5115 0.6405
4.1.2 Replica
Table 4: Video Object Segmentation results on the Replica[50] dataset. Positive IoU is computed over frames with ground-truth masks. Successful IoU is computed over frames where a ground-truth object exists and the IoU of prediction and GT $\neq 0$.
Method IoU $\uparrow$ Positive IoU $\uparrow$ Successful IoU $\uparrow$

### v5

@@ Abstract
Video object segmentation methods like SAM2 achieve strong performance through memory-based architectures but struggle under large viewpoint changes due to reliance on appearance features. Traditional 3D instance segmentation methods address viewpoint consistency but require camera poses, depth maps, and expensive preprocessing. We introduce 3AM, a training-time enhancement that integrates 3D-aware features from MUSt3R into SAM2. Our lightweight Feature Merger fuses multi-level MUSt3R features that encode implicit geometric correspondence. Combined with SAM2’s appearance features, the model achieves geometry-consistent recognition grounded in both spatial position and visual similarity. We propose a field-of-view aware sampling strategy ensuring frames observe spatially consistent object regions for reliable 3D correspondence learning. Critically, our method requires only RGB input at inference, with no camera poses or preprocessing. On challenging datasets with wide-baseline motion (ScanNet++, Replica), 3AM substantially outperforms SAM2 and extensions, achieving 90.6% IoU and 71.7% Tracking Recall on ScanNet++’s Selected Subset, improving over state-of-the-art VOS methods by +15.9 and +30.4 points. Project page: https://jayisaking.github.io/3AM-Page/
@@ 1 Introduction
Both paradigms have critical gaps (Fig.˜2). 2D methods like SAM2 excel at efficiency but lack geometric awareness, failing under viewpoint variation (Fig.˜2a). MOSEv2 [17] shows significant degradation under wide-baseline changes. Conversely, 3D methods achieve better consistency but require explicit 3D inputs (Fig.˜2b) and struggle to generalize. 2D-to-3D lifting suffers from view-inconsistent predictions [31, 58, 73, 94, 116]. End-to-end approaches like PanSt3R [123] require offline processing. Can we achieve 3D-aware, viewpoint-robust segmentation while preserving promptability and efficiency, without explicit 3D supervision? (Fig.˜2c)
We introduce 3AM (Fig.˜1), a generalized 3D-aware tracker. Our key insight is that MUSt3R [7] encodes implicit geometric correspondence through features learned from multi-view consistency. By fusing multi-level MUSt3R features with SAM2’s appearance features via a lightweight Feature Merger, we enable geometry-consistent recognition based on spatial position, which at inference requires only RGB input and user prompts. We introduce a field-of-view aware sampling strategy ensuring frames observe spatially consistent object regions during training. On ScanNet++ [109] and Replica [69] with wide-baseline motion, 3AM achieves 90.6% IoU and 71.7% Positive IoU versus SAM2Long’s 74.7% and 41.3% (+15.9 and +30.4 points).
@@
  • A generalized tracking framework that adaptively reverts to the original SAM2 for non-rigid or highly dynamic objects, along with extensive experimental evaluation demonstrating consistent improvements over state-of-the-art VOS methods on challenging datasets such as ScanNet++ and Replica, especially under object reappearance and significant viewpoint changes, while retaining competitive performance on conventional VOS benchmarks.
@@ 2 Related Work
2.0.1 2D Video Object Segmentation.
Video object segmentation (VOS) segments and tracks target objects across video sequences with temporally coherent masks [59, 8, 78, 51, 87]. Early approaches relied on appearance propagation or optical flow but struggled with long-term consistency and occlusion [100, 55, 4]. Recent memory-based architectures leverage spatio-temporal attention and feature retrieval for improved object association [104, 13, 14, 47, 56, 93, 103, 105, 42, 12], with subsequent work improving memory efficiency via restricted banks [118], gated linear matching [46], unified transformer frameworks [43, 114], and dynamic anchor queries [120]. New benchmarks [26] and specialized settings such as panoramic video [97] and point-based prompting [49] further broaden the evaluation landscape.
SAM2 [63], a promptable model with streaming memory, achieves strong results across benchmarks. Follow-up efforts extend SAM2 with memory trees for long-range identity preservation [18], distractor-aware updates [77], on-device efficiency [117], text-driven understanding [15], and other improvements [98, 101, 108]. Concurrent work further combines SAM2 with large language models for reasoning-based segmentation [112, 96, 2]. Despite these advances, large viewpoint changes, occlusion, and disappearance-reappearance events cause significant degradation [17, 92, 61]. Our method addresses these limitations by integrating 3D-aware representations into SAM2 for object-consistent segmentation under challenging viewpoint and appearance variations.
2.0.2 3D Instance Segmentation.
Mainstream 3D instance segmentation follows two tracks. Proposal-centric methods predict instance masks from point clouds with learned proposals [65, 34, 25, 80, 52, 71, 38, 73, 60], with recent extensions adding promptable segmentation [121] and fast open-vocabulary recognition [6]. A complementary line lifts 2D results to 3D by projecting per-view masks and merging across views [102, 94, 54, 5, 35, 90, 36, 79, 33, 3, 27, 76, 41], employing superpoint affinity [110], zero-shot feature distillation [84, 20], hierarchical contrastive learning [111, 37], and egocentric Gaussian lifting [24], but still depending on posed RGB-D for mask consolidation.
The rise of 3D Gaussian Splatting has further enabled semantic scene representations by distilling foundation-model features into Gaussians [62, 119, 88, 57], assigning per-Gaussian identity encodings [107, 113], solving 2D-to-3D label assignment optimally [66], and unifying lifting with segmentation end-to-end [122]. However, pure 2D promptable segmentation rarely suffices for 3D instances because cross-view consistency is not enforced, leading to view-inconsistent masks and fragmented instances [116, 94, 58, 73, 31]. Our approach instead imposes 3D-aware self-consistency directly on video, preserving identity and geometry without 3D annotations or point-cloud mask merging.
@@
2.0.3 End-to-End 3D-Aware Methods.
Recent end-to-end 3D reconstruction models directly infer geometric structure from 2D inputs [7, 86, 85, 83, 99, 11, 39, 89, 9, 106, 81, 44, 67, 21, 48, 70]. The DUSt3R/MASt3R family has been extended to dynamic scenes [115], incremental reconstruction with spatial memory [82], feed-forward Gaussian splatting [68], efficient multi-view fusion [74], real-time dense SLAM [53], unconstrained structure-from-motion [19], and training-free 4D reconstruction [10]. Feed-forward models now also jointly predict geometry and semantics [23, 72, 32], with studies analyzing how different foundation models contribute complementary 3D capabilities [50].
This progress has fueled end-to-end 3D-aware segmentation frameworks [29, 30, 40]. PanSt3R [123] integrates 3D-aware features from MUSt3R [7] with a transformer decoder for joint geometry and panoptic segmentation via learnable quadratic fusion. However, it is incompatible with promptable backbones (e.g., SAM2), limiting interactivity, and requires offline access to entire sequences, preventing streaming.
@@ 3.1 Problem Setting
Given an RGB image $\mathbf{I}$ and a user-provided prompt $p$ (mask, box, or point), our goal is to identify the same object throughout a video sequence with consistent segmentation, avoiding false correspondences or identity switches across frames. Previous approaches often fail at this goal. 2D-based methods (e.g., SAM2 [63]) rely solely on appearance features, making them sensitive to viewpoint or distance changes and prone to confusion among visually similar objects [77], leading to frequent tracking failures over longer sequences. 3D-based methods (e.g., Mask3D [65], OpenYOLO3D [5], EmbodiedSAM [95]) localize objects using 3D mask proposals or dense 2D masks with camera poses, but require SfM or similar pipelines whose runtime and complexity grow rapidly with the number of frames, making them impractical for long sequences. One might instead apply a 3D segmentation model on predicted geometry (e.g., point clouds from MUSt3R). However, geometry quality, especially in dynamic scenes, limits downstream segmentation reliability. Moreover, current 3D segmentation models are mainly offline, unsuitable for real-time scenarios such as robotic navigation. More importantly, large-scale training data is far more abundant in the 2D video domain, and the SAM2 model we adopt provides a strong, promptable 2D prior difficult to match with current 3D alternatives. Our method leverages this advantage by using 3D information only as a lightweight association cue rather than a hard prerequisite, maintaining a 2D segmentation framework. Specifically, we encourage geometry-consistent feature matching, enabling the model to identify the same object by recognizing its correspondence in 3D space.
@@ 3.2 Architecture
The overall pipeline is shown in Fig.˜3. For each video frame $i$, our model extracts two complementary feature streams that capture appearance cues and multi-view geometric consistency. First, the frame is processed by the SAM2 vision encoder, producing a 2D appearance feature map $F_{\mathrm{2D}}^{i}$. In parallel, the same frame is passed into MUSt3R, which attends to its own multi-view memory through MUSt3R’s internal cross-attention mechanism, yielding a geometry-aware feature map $F_{{3D}}^{i}$. Next, $F_{{2D}}^{i}$ and $F_{{3D}}^{i}$ are fused into a refined representation $F_{{merged}}^{i}$. The merged feature is processed by the memory attention module and the mask decoder, and is finally encoded by the memory encoder to produce memory features that will be referenced by future frames. We provide comparisons of different memory selection methods in Tab.˜8.
3.2.1 Feature Merging.
@@
As shown in Fig.˜4, SAM2 becomes unreliable when the object reappears or when the viewpoint changes drastically. The heat map in the lower row demonstrates that SAM2 cannot reliably associate the query point with its target under large angular differences, revealing its limited capacity to maintain object identity when the visual appearance changes substantially.
To address this issue, we enhance SAM2 with MUSt3R features $F_{{3D}}^{i}$. The upper row of Fig.˜4 visualizes MUSt3R activations across depth. Early layers preserve a clearer correspondence pattern with the query point, while deeper layers produce more diffuse responses. This shift reflects MUSt3R’s progression toward point-cloud decoding: the representation becomes more geometry-oriented but less semantically aligned with the 2D query. Consequently, relying solely on either early or late layers is insufficient; early layers lack explicit geometric structure, whereas late layers lose semantic clarity. We therefore sample MUSt3R features from both early and late stages to capture complementary information.
These multi-level MUSt3R features are processed by our Feature Merger, which jointly performs cross-attention fusion and convolutional refinement. The cross-attention blocks first integrate the sampled MUSt3R layers into a single geometry-aware feature, aligning information from different depths. Specifically, as shown in Fig.˜4, we take the MUSt3R encoder feature at the shallowest sampled depth, i.e., the most semantically rich layer, as the initial input to the Feature Merger. It first passes through a self-attention block to establish an initial semantic representation. The remaining sampled MUSt3R features are then incorporated one by one through a sequence of cross-attention layers, where the current merged representation acts as the query, and each additional MUSt3R feature provides the key–value pairs. In this way, the Feature Merger progressively integrates information from both shallow and deeper stages of MUSt3R. The output is then fed into a convolutional stage that merges it with SAM2’s 2D feature $F_{2D}$ for restoring fine spatial detail. We further compare different 3D foundation models in Tab.˜8, and more architectural details and comparisons are provided in the supplementary materials.
The final merged feature, $F_{{merged}}$, retains SAM2’s high-resolution, segmentation-aware appearance cues while incorporating MUSt3R’s geometry-aware consistency, enabling robust object localization even under severe viewpoint changes.
@@ 3.3 Training Frame Sampling
3.3.1 Naïve Random Sampling.
A straightforward strategy is to randomly sample $N$ frames from the video as in Fig.˜5(b). This exposes the model to a broad distribution of camera viewpoints and object appearances. Random sampling also naturally introduces distracting or visually similar objects in the scene , which encourages the model to rely on 3D-aware alignment rather than purely local 2D similarity.
3.3.2 Object Spanning Problem.
@@
3.3.3 Field-of-View Sampling.
@@
Note that we apply FoV Sampling as a filter only when candidate masks contain an object. For no-object frames, which are also sampled during training, the model is simply required to predict target absence. In these cases, spatial distance does not affect the objective. That is, when views are far apart, the prediction is driven by geometric inconsistency, while for nearby views, it relies on appearance cues.
We describe our final sampling policy and its ablations in Sec.˜4.
3.4 Dynamic Object Fallback
An intuitive limitation of our method lies in dynamic objects, whose shape, location, or orientation may change over time. Such scenarios are prevalent in traditional VOS benchmarks [61, 92, 22]. Our approach relies heavily on relative appearance with respect to surrounding geometry (i.e., consistent geometrical locations across views). When an object undergoes significant motion, this assumption may break, potentially leading to tracking failure. To address this issue, we detect the presence of motion using [28] and revert to the original SAM2 pipeline when substantial motion is observed. Since the SAM2 image encoder remains frozen in our framework, we can reuse the visual feature and the additional computational overhead incurred by memory attention and mask decoder is negligible.
Empirically, we observed that dynamically toggling between 3AM and vanilla SAM2 mid-sequence (e.g., with a simple motion detection branch before mask decoder) corrupts the temporal memory bank, as the two pathways produce fundamentally different mask distributions. Therefore, we utilize a simple sequence-level fallback to the original SAM2 pipeline when substantial motion is detected. This design choice is grounded in a practical categorization of tracking scenarios based on the object’s visibility with respect to the camera frustum: (1) Dynamic object with continuous visibility, where the object remains within the camera frustum and vanilla SAM2 performs well; (2) Static object with temporary disappearance, where the object leaves the camera frustum and later re-enters, resulting in a re-identification problem that 3AM is designed to address; and (3) Dynamic object with unobserved motion outside the frustum.
The third scenario represents an inherently ill-posed problem, as the object may move unpredictably while unobserved outside the camera frustum. Consequently, reliable tracking is fundamentally impossible. Since this extreme scenario cannot be effectively solved, engineering a complex mid-sequence fusion module is unnecessary. Hence a simple sequence-level switch is sufficient to cover the two tractable scenarios.
@@ 4 Experiments
To mitigate the imperfect 2D mask annotations common in 3D-based datasets, we train on a hybrid mixture of synthetic, real, and video segmentation datasets. Specifically, the Aria Synthetic Dataset(ASE) [1] provides clean geometric supervision, ScanNet++ [109] offers realistic indoor 3D consistency, and MOSE[16] maintains the original temporally coherent masks. This combination enables balanced spatial and temporal learning. Note that to remain consistent with our sampling strategy described in Sec.˜3.3, we apply continuous sampling on MOSE, as it does not provide camera poses or depth maps. For ASE and ScanNet++, which include calibrated geometry, we adopt a mixed policy: continuous sampling with probability 0.2 and our FOV-aware sampling with probability 0.8. The FOV threshold $\tau$ is set to 0.25. To apply our sampling strategy, we first run MUSt3R inference on every scene in ScanNet++ and MOSE and store the resulting features. These precomputed features are then used during training to enable FOV-aware sampling. For continuous sampled batches, MUSt3R is executed on the fly.
@@
To address this limitation, we use 3D-based datasets, ScanNet++ [109] and Replica [69], which naturally provide extensive camera trajectory variation due to their 3D reconstruction requirements. Their wide-baseline viewpoints make them suitable for evaluating consistency under severe pose changes. For each scene in both datasets, we use ground-truth camera poses to project 3D instance masks onto the 2D frames. We then select the frame with the largest visible mask area as the conditioning frame and perform forward and backward tracking from that reference.
4.1.1 Compared Methods.
For comparison, we use SAM2 as the state-of-the-art VOS baseline. We also include the following works: SAM2Long [18] and DAM4SAM [77], which enhance SAM2 by improving its memory selection for more robust tracking. Note that our method, 3AM, uses the naive memory mechanism of the original SAM2 in all reported results unless specified otherwise.
4.1.2 Metrics.
We use three complementary metrics to evaluate tracking quality. IoU is computed over all frames, including those where the object is absent, reflecting overall stability. Tracking Recall measures accuracy only on frames where a target object is present. Accuracy is further restricted to visible frames where the prediction has non-zero overlap with ground truth, capturing accuracy conditioned on successful localization. These metrics assess robustness to disappearance and accuracy when tracking succeeds.
4.1.3 ScanNet++.
Table 2: Video Object Segmentation results on the ScanNet++ [109] and VOS datasets. Tracking Recall is computed over frames with object presence. Accuracy is computed over frames where a ground-truth object exists and the IoU of prediction and GT $\neq 0$.
Method FPS Mem ScanNet++ Whole Set ScanNet++ Selected Subset $\text{DAVIS}_{17}$ LaSOT $\text{LaSOT}_{\text{ext}}$ DiDi
G IoU $\uparrow$ Tracking Recall $\uparrow$ Accuracy $\uparrow$ IoU $\uparrow$ Tracking Recall $\uparrow$ Accuracy $\uparrow$ J&F $\uparrow$ AUC $\uparrow$ AUC $\uparrow$ Accuracy $\uparrow$
SAM2 [63] 14.4 3.6 0.4392 0.0235 0.0831 0.3397 0.0179 0.0395 89.0 70.0 56.9 0.720
[…]
@@
The results are shown in Tab.˜2, and we provide visualization at Fig.˜8. On the ScanNet++ dataset, we further evaluate performance on a Selected Subset specifically constructed to emphasize object reappearance. For each object track, we count the number of non-continuous visible segments, where a segment is defined as any contiguous span of visible frames whose length $\ell$ exceeds a minimal threshold $\ell_{\min}$. Objects with frequent reappearance are typically associated with large viewpoint changes, since each disappearance–reappearance cycle often corresponds to observing the object from a substantially different angle or position. This subset, therefore, provides a focused evaluation of robustness under severe pose variation.
3AM achieves the best performance on both the Whole Set (IoU 0.8898, highest Recall and Accuracy) and the challenging Selected Subset (IoU 0.9061, Recall 0.7168, Accuracy 0.7737), highlighting its ability to maintain object identity even when the object reappears under large changes in viewpoint and spatial configuration. With a simple fallback mechanism, it also retains competitive results on conventional VOS benchmarks [61, 22, 77] as shown in Tab.˜2.
4.1.4 Two-view Matching Comparison with SegMASt3R.
We further compare our 3AM with SegMASt3R [30], a two-view matching approach built on MASt3R that associates segment masks between two images under large viewpoint changes, in Tab.˜4. SegMASt3R is not directly prompable and does not natively support >2>2 multiview tracking. Nevertheless, we include it as a strong two-view baseline by adopting the following protocol: we use the ground-truth initial mask as the source-frame candidate, and for each subsequent frame, we independently perform two-view mask matching against the source frame. To fairly compare the two approaches, when tracking with our 3AM, we don’t use any memory other than the source frame.
Under this protocol, 3AM achieves IoU of $0.8915$, Tracking Recall of $0.5115$, and Accuracy of $0.6405$, whereas SegMASt3R obtains $0.6800$, $0.3628$, and $0.4053$, respectively. These results demonstrate 3AM not only provides stronger functionality, but also outperforms SegMASt3R even when restricted to two-view matching.
Table 3: Quantitative comparison of two-view matching between 3AM and SegMASt3R on ScanNet++ Whole Set.
Method IoU $\uparrow$ Tracking Recall $\uparrow$ Accuracy $\uparrow$
@@
[…]
Table 4: Video Object Segmentation results on the Replica[69] dataset.
Method IoU $\uparrow$ Tracking Recall $\uparrow$ Accuracy $\uparrow$